Ammonium nitrate (NH.NO,) decomposes to yield nitrogen gas, water, and oxygen gas.


a. Write the balanced chemical reaction:


b. How many grams of water are produced when 36.0 grams of ammonium

nitrate decompose?



1
Expert's answer
2021-05-25T07:51:22-0400

Solution:

Ammonium nitrate = NH4NO3

a) Balanced chemical equation:

2NH4NO3(s) → 2N2(g) + O2(g) + 4H2O(g)


b) The molar mass of NH4NO3 is 80.0 g/mol.

Therefore,

(36.0 g NH4NO3) × (1 mol NH4NO3 / 80.0 g NH4NO3) = 0.45 mol NH4NO3


According to the equation above: n(NH4NO3) = n(H2O)/2

n(H2O) = 2 × n(NH4NO3) = 2 × 0.45 mol = 0.90 mol


The molar mass of H2O is 18.0 g/mol.

Therefore,

(0.90 mol H2O) × (18.0 g H2O / 1 mol H2O) = 16.2 g H2O


Answers:

a) 2NH4NO3(s) → 2N2(g) + O2(g) + 4H2O(g)

b) 16.2 g H2O